a simple random sample of size 5 is drawn without replacement from a finite population consisting of 41 units. if the population standard deviation is 6.25 . what is the standard error of sample mean?


1
Expert's answer
2021-02-24T06:35:43-0500
"\sigma_{\bar{x}}=\sqrt{\dfrac{N-n}{Nn}}\sigma"

"\sigma_{\bar{x}}=\sqrt{\dfrac{41-5}{41(5)}}\cdot6.25=2.62"